Today you sell your stock fund for $32,092. You bought it 5 years ago and otherwise the account has been left alone. The stocks have earned a 13% annual rate of return. How much did you buy the stock for?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

32092=P*(1.13)^5

P=32092/(1.13)^5

=32092*0.542759936

which is equal to

=$17418.25(Approx).